OpenStudy (anonymous):

Find the x-intercepts of f(x)=3x^2-5x-11=0

OpenStudy (anonymous):

maybe use quad. formula or complete square

OpenStudy (ybarrap):

http://www.wolframalpha.com/input/?i=plot+3x^2-5x-11%3D0 3 x^2-5 x-11 = 0 Solve the quadratic equation by completing the square: Divide both sides by 3: x^2-(5 x)/3-11/3 = 0 Add 11/3 to both sides: x^2-(5 x)/3 = 11/3 Add 25/36 to both sides: x^2-(5 x)/3+25/36 = 157/36 Factor the left hand side: (x-5/6)^2 = 157/36 Take the square root of both sides: abs(x-5/6) = sqrt(157)/6 Eliminate the absolute value: x-5/6 = -sqrt(157)/6 or x-5/6 = sqrt(157)/6 Add 5/6 to both sides: x = 1/6 (5-sqrt(157)) or x-5/6 = sqrt(157)/6 Add 5/6 to both sides: x = 1/6 (5-sqrt(157)) or x = 1/6 (5+sqrt(157)) Ans. x = 1/6 (5 +/- sqrt(157))